# Distributionally Robust Policy Learning under Concept Drifts

### v1

@@ 4 Policy learning under concept drift
$\mathbb{E}\big{[}\ell(X,Y(\pi(X));\theta)\,|\,X=x\big{]}=\sum^{M}_{a=1}\mathbbm{1}\{\pi(X)=a\}\cdot\mathbb{E}[\ell(x,Y(a);\theta)\,|\,X=x].$ (8)
@@
$-\mathbb{E}\Bigg{[}\bm{\alpha}^{*}_{\pi(X)}(X)\cdot\exp\bigg{(}-\frac{Y(\pi(X))+\bm{\eta}^{*}_{\pi(X)}(X)}{\bm{\alpha}^{*}_{\pi(X)}(X)}-1\bigg{)}+\bm{\eta}^{*}_{\pi(X)}(X)+\bm{\alpha}^{*}_{\pi(X)}(X)\delta\Bigg{]}.$
@@
Constructing $\widehat{\bm{\theta}}_{\pi}(x)$ with $\widehat{\bm{\theta}}_{\pi(x)}(x)$ substantially reduces the computational complexity of the policy learning task. It is virtually infeasible to estimate $\widehat{\bm{\theta}}_{\pi}(x)$ for each $\pi$ in a policy class $\Pi$ with infinite number of policies. Alternatively, solving for $\widehat{\bm{\theta}}_{\pi(x)}(x)$ transforms the infeasible task of computing a class of infinite nuisance parameters $\{\bm{\theta}_{\pi}(x):\pi\in\Pi\}$ to the feasible task of computing a finite one $\{\bm{\theta}_{a}(x):a\in[M]\}$. It remains an interesting future direction to extend this trick to account for continuous action spaces.
The following section instantiates this idea and provides a detailed policy learning algorithm.
@@ 4.1 The learning algorithm
The policy learning algorithm consists of two main steps:
  • (1)
    solving for $\bm{\theta}^{*}_{a}$ for each $a\in[M]$ and constructing the policy value estimator $\widehat{\mathcal{V}}_{\delta}(\pi)$;
  • (2)
    learning the optimal policy $\pi^{*}_{\delta}$ by minimizing $\widehat{\mathcal{V}}_{\delta}(\pi)$.
@@
$G_{a}(x,y)=\ell(x,y;\bm{\theta}^{*}_{a}(x)),~{}\widehat{G}^{(k)}_{a}(x,y)=\ell(x,y;\widehat{\bm{\theta}}^{(k)}_{a}(x)),\text{ and }\bar{g}^{(k)}_{a}(x)=\mathbb{E}\big{[}\widehat{G}^{(k)}_{a}(X,Y(a))\,|\,X=x\big{]}.$
@@
$\widehat{\pi}_{\text{LN}}=\underset{\pi\in\Pi}{\arg\!\max}~{}\widehat{\mathcal{V}}^{\textnormal{LN}}_{\delta}(\pi):=-\frac{1}{K}\sum^{K}_{k=1}\widehat{\mathcal{V}}_{\delta}^{\text{LN},(k)}(\pi),\text{ where }$ (9)
$\widehat{\mathcal{V}}_{\delta}^{\text{LN},(k)}(\pi)=\frac{1}{|\mathcal{D}^{(k)}|}\sum_{i\in\mathcal{D}^{(k)}}\frac{\mathbbm{1}\{A_{i}=\pi(X_{i})\}}{\widehat{\pi}_{0}^{(k)}(A_{i}\,|\,X_{i})}\cdot\big{(}\widehat{G}^{(k)}_{\pi(X_{i})}(X_{i},Y_{i})-\widehat{g}^{(k)}_{\pi(X_{i})}(X_{i})\big{)}+\widehat{g}^{(k)}_{\pi(X_{i})}(X_{i}).$
The above optimization problem can be solved efficiently by first-order optimization methods or policy tree search as in Zhou et al., (2023); we shall elaborate on the implementation in Section 5. The complete policy learning procedure is summarized in Algorithm 2.
@@
Return: $\widehat{\pi}_{\text{LN}}$ that maximizes $\widehat{\mathcal{V}}_{\delta}^{\text{LN}}(\pi)$ as in Equation (LABEL:eqn:policyvaluelearn).
@@ 4.2 Regret upper bound
Definition 4.2 (Hamming entropy intergal).
@@
The proof of Theorem 4.3 is deferred to Appendix B. At a high level, we decompose the regret and upper bound it with the supremum of the estimation error of policy values:
$\mathcal{R}_{\delta}(\widehat{\pi}_{\text{LN}})$ $=\mathcal{V}_{\delta}(\pi^{\ast})-\mathcal{V}_{\delta}(\widehat{\pi}_{\text{LN}})\leq\mathcal{V}_{\delta}(\pi^{*})-\widehat{\mathcal{V}}_{\delta}^{\text{LN}}(\pi^{*})+\widehat{\mathcal{V}}_{\delta}^{\text{LN}}(\pi^{*})-\widehat{\mathcal{V}}_{\delta}^{\text{LN}}(\widehat{\pi}_{\text{LN}})+\widehat{\mathcal{V}}_{\delta}^{\text{LN}}(\widehat{\pi}_{\text{LN}})-\mathcal{V}_{\delta}(\widehat{\pi}_{\text{LN}})$
$\leq 2\sup_{\pi\in\Pi}|\widehat{\mathcal{V}}^{\textnormal{LN}}_{\delta}(\pi)-\mathcal{V}_{\delta}(\pi)|,$
where the last step uses the choice of $\widehat{\pi}_{\textnormal{LN}}$. We bound the above quantity by establishing uniform convergence results for the policy value estimators. Through a careful chaining argument, we show that the dependence of $\mathcal{R}(\widehat{\pi}_{\text{LN}})$ on $n$ is of the order $O(n^{-\frac{1}{2}})$, which is sharper than the $O(n^{-\frac{1}{2}}\log n)$ dependence for that of  Mu et al., (2022) by a logarithmic factor. We also note that both regrets are asymptotic in $n$ and hold for sufficiently large $n$.
@@ 4.3 Regret lower bound
In this section, we complement the regret upper bound in Theorem 4.3 with a minimax lower bound that characterizes the fundamental difficulty of policy learning under concept drift. Our lower bound is stated in terms of the Natarajan dimension (Natarajan,, 1989), defined as follows.
@@
Given an $M$-action policy class $\Pi$, we say a set of $m$ points $\{x_{1},\ldots,x_{m}\}$ is shattered by $\Pi$ if there exist two functions $f_{-1},f_{1}:\{x_{1},\ldots,x_{m}\}\mapsto[M]$ such that
  • (1)
    $f_{-1}(x_{j})\neq f_{1}(x_{j})$ for all $j\in[m]$;
  • (2)
    for any $\sigma\in\{\pm 1\}^{m}$, there exists a policy $\pi\in\Pi$ such that for any $j\in[m]$, $\pi(x_{j})=f_{\sigma_{j}}(x_{j})$.
We define the Natarajan dimension Ndim($\Pi$) of $\Pi$ to be the size of the largest set shattered by $\Pi$.
@@
As can be seen from the definition, the Natarajan dimension generalizes the Vapnik-Chervonenkis (VC) dimension (Vapnik and Chervonenkis,, 2015) to the multi-class classification setting. The Natarajan dimension is also closely related to the Hamming entropy integral $\kappa(\Pi)$ in our upper bound, as $\kappa(\Pi)=O(\sqrt{\log(d)\text{Ndim}(\Pi)}$(Cai et al.,, 2020).
@@
Let $\mathcal{P}$ denote the set of all distributions of $(X,A,Y(1),\ldots,Y(M))$ that satisfy Assumption 2.1, 2.4, 3.3, and 3.4.555When we say a distribution $P$ satisfies Assumption 3.4, we mean that under $P$ there exist $\widehat{\bm{\theta}}$, $\widehat{\pi}_{0}$, and $\widehat{g}$ that satisfy the convergence rates in Assumption 3.4. Suppose that $\delta\leq 0.2$, $n\geq\textnormal{Ndim}(\Pi)^{2}$, and $\textnormal{Ndim}(\Pi)\geq 4/(9\varepsilon)$. For any policy leaning algorithm that outputs $\widehat{\pi}$ as a function of $\{(X_{i},A_{i},Y_{i})\}_{i=1}^{n}$, there is
$\sup_{P\in\mathcal{P}}\mathbb{E}_{P^{n}}[\mathcal{R}(\widehat{\pi})]\geq\frac{1}{120}\sqrt{\frac{\textnormal{Ndim}(\Pi)}{n\varepsilon}}.$
The proof of Theorem 4.6 is provided in Appendix B.4. Theorem 4.6 implies that for any learning algorithm, there exists a problem instance such that the regret scales as $\Omega(\sqrt{\text{Ndim}(\Pi)/n})$.
@@
Recalling the relationship between the Natarajan dimension and the Hamming entropy integral in the remark above, we see that our proposed algorithm achieves the minimax rate in the sample size and the policy class complexity up to logarithmic factors.
5 Numerical results
We evaluated our learning algorithm in a simulated setting against the benchmark algorithm SNLN in Si et al., (2023).
Data generating process.
Our data generating process follows that of the linear boundary example in Si et al., (2023). We let the context set $\mathcal{X}=\{x\in\mathbb{R}^{5}:\|x\|_{2}\leq 1\}$ to be the closed unit ball of $\mathbb{R}^{5}$ and let the action set to be $\mathcal{A}=\{1,2,3\}$; the rewards $Y(a)$’s are mutually independent conditioned on $X$ with $Y(a)\mid X\sim N(\beta_{a}^{\top}X,\sigma_{a}^{2})$, for $a\in[3]$. The training dataset $\mathcal{D}_{\text{train}}=\{(X_{i},Y_{i}(\pi_{0}(X_{i})))\}_{i=1}^{n}$ are generated with a given behavior policy $\pi_{0}$ (unknown to policy learning algorithms). The testing datasets $\mathcal{D}_{\text{test}}$ are i.i.d. draws of data tuple $\{(X_{i},Y_{i}(1),Y_{i}(2),Y_{i}(3))\}_{i=1}^{n}$. The specific values of $\{\beta_{1},\beta_{2},\beta_{3}\}\in\mathbb{R}^{5}$ and $\{\sigma^{2}_{1},\sigma^{2}_{2},\sigma^{2}_{3}\}\in\mathbb{R}_{+}$ as well as the behavior policy $\pi_{0}$ are given in Appendix A.
Implementation.
In our implementation, the number of splits is taken to be $K=3$. We use the Random Forest regressor from the scikit-learn Python library to estimate $\widehat{\pi}_{0}$ and $\widehat{g}$. For estimating $\bm{\theta}^{*}$, we adopt the cubic spline method and employ the Nelder-Mead optimization method in SciPy Python library (Virtanen et al.,, 2020) to optimize the coefficients in the spline approximation, where the obtained estimator has threshold at 0.001 to guarantee Proposition 2.5. Finally, we optimize and find $\widehat{\pi}_{\textnormal{LN}}$ with policytree (Athey and Wager,, 2021).
The benchmark algorithm SNLN is adapted from Si et al., (2023, Algorithm 2). Since Si et al., (2023, Algorithm 2) is designed for joint distribution shift formulation, we revised the original algorithm to fit our concept drift setting. It is well-known that the chain rule of KL-divergence (Cover,, 1999) gives
$D_{\text{KL}}(Q_{X,Y}\,\|\,P_{X,Y})=D_{\text{KL}}(Q_{X}\,\|\,P_{X})+D_{\text{KL}}(Q_{Y\,|\,X}\,\|\,P_{Y\,|\,X}).$ (10)
Therefore, given any uncertainty set radius $\delta$ and known covariate shift (in this experiment, we assume no covariate shift), Si et al., (2023, Algorithm 2) can be used to implement policy learning under concept drift. Note that SNLN admits known propensity scores. As we only consider the case where the propensity scores are unknown, we complement Si et al., (2023, Algorithm 2) with estimated propensity scores from Random Forest Regressor in scikit-learn, the same way as in the implementation of Algorithm 2. The additional setup details are in Appendix A.
Evaluation.
For a learned policy $\widehat{\pi}$, we evaluate its performance by the following two performance metrics. (i) We use the testing dataset $\mathcal{D}_{\text{test}}$ to estimate the robust policy value $\mathcal{V}^{*}_{\delta}(\widehat{\pi})$ by the empirical robust policy value
$\bar{\mathcal{V}}_{\delta}(\widehat{\pi})=-\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{i\in\mathcal{D}_{\text{test}}}\bigg{[}\bm{\alpha}_{\widehat{\pi}}(X_{i})\exp\Big{(}-\frac{Y_{i}(\widehat{\pi}(X_{i}))+\bm{\eta}_{\widehat{\pi}}(X_{i})}{\bm{\alpha}_{\widehat{\pi}}(X_{i})}-1\Big{)}+\bm{\eta}_{\widehat{\pi}}(X_{i})+\bm{\alpha}_{\widehat{\pi}}(X_{i})\delta\bigg{]},$
where the nuisance parameters $\bm{\alpha}_{\widehat{\pi}}(X_{i}),\bm{\eta}_{\widehat{\pi}}(X_{i})$ are found via cubic spline method and employ the Nelder-Mead optimization method using the testing dataset according to policy $\widehat{\pi}$: $\mathcal{D}_{\text{test},\widehat{\pi}}=\{(X_{i},Y_{i}(\widehat{\pi}(X_{i})))\}$. (ii) For 100 testing dataset each containing 10000 data points $\big{\{}\{(X_{i}^{(j)},Y_{i}^{(j)}(1),Y_{i}^{(j)}(2),Y_{i}^{(j)}(3))\}_{i=1}^{10000}\big{\}}_{j=1}^{100}$, we randomly sample a new testing dataset $\big{\{}\{(X_{i}^{(j)},\tilde{Y}_{i}^{(j)}(1),\tilde{Y}_{i}^{(j)}(2),\tilde{Y}_{i}^{(j)}(3))\}_{i=1}^{10000}\big{\}}_{j=1}^{100}$ on the KL-sphere centered at $(Y_{i}^{(j)}(1),Y_{i}^{(j)}(2),Y_{i}^{(j)}(3))$ with a radius $\delta$. Then we evaluate $\widehat{\pi}$ using
$\tilde{\mathcal{V}}_{\delta}^{\min}(\widehat{\pi}):=\min_{1\leq j\leq 100}\bigg{\{}\frac{1}{10000}\sum_{i=1}^{10000}\tilde{Y}_{i}^{(j)}\big{(}\widehat{\pi}(X_{i}^{(j)})\big{)}\bigg{\}}.$
Results.
Table 2 and 3 report the values $\bar{\mathcal{V}}_{\delta},\tilde{\mathcal{V}}_{\delta}^{\min}$ (with 95% confidence intervals) of the learnt policies $\widehat{\pi}_{\text{LN}}$ and $\widehat{\pi}_{\text{SNLN}}$, by Algorithm 2 and Si et al., (2023, Algorithm 2) respectively. Table 2 shows that $\widehat{\pi}_{\text{LN}}$ outperforms the benchmark $\widehat{\pi}_{\text{SNLN}}$ consistently, with higher policy values and similar 95% confidence intervals. With a higher $\delta$, the policy values of $\widehat{\pi}_{\text{LN}},\widehat{\pi}_{\text{SNLN}}$ are smaller, due to a bigger uncertainty set. Table 3 shows that $\widehat{\pi}_{\text{LN}}$ achieves higher worst-case rewards than $\widehat{\pi}_{\text{SNLN}}$ does, in a more realistic setting with concept drift testing datasets. Together, we see that $\widehat{\pi}_{\text{LN}}$ succeeds in finding a better policy under concept drift; while the performance of $\widehat{\pi}_{\text{SNLN}}$ is comprised by its conservative policy learning process, in which it considers joint distributional shifts even though it is given the information that no covariate shifts took place.
The results align with the intuition that Algorithm 2 admits a subset of the uncertainty set that the benchmark algorithm SNLN considers, as explained in Equation (10). Consequently, $\mathcal{V}_{\delta}(\widehat{\pi}_{\text{SNLN}})$ is a lower bound of $\mathcal{V}_{\delta}(\widehat{\pi}_{\text{LN}})$ in theory, and by the results in Table 2, in practice. In real-world applications, knowing the source of the distribution shift effectively shrinks the uncertainty set, thereby yielding less conservative results. Since it is fairly easy to identify covariate shifts (comparing to detecting concept drift), when the decision maker observes none or little covariate shifts and would like to hedge against the risk of concept drift, it is suitable to apply our method which outperforms existing method designed for learning under the joint distributional shift.
In Appendix A, we also provide simulation results of Algorithm 1 for a fixed target policy, which show that Algorithm 1 can estimate the distributionally robust policy value under concept drift efficiently.
$n=$7500 $n=$13500 $n=$16500 $n=$19500
$\bar{\mathcal{V}}_{0.05}$ $\widehat{\pi}_{\text{LN}}$ 0.2272±$\pm 0.002$ 0.2299±$\pm 0.001$ 0.2303±$\pm 0.001$ 0.2310±$\pm 0.001$
@@ Results.
Table 2: Empirical robust policy value $\bar{\mathcal{V}}_{\delta}$ of policies $\widehat{\pi}_{\text{LN}},\widehat{\pi}_{\text{SNLN}}$ learned by Algorithm 2 and SNLN respectively, under $\delta=0.05,0.1,0.2$, each with 50 trials.
$n=$7500 $n=$13500 $n=$16500 $n=$19500
@@
Table 3: Empirical worst case policy reward on the KL-sphere $\tilde{\mathcal{V}}_{\delta}^{\min}$ of policies $\widehat{\pi}_{\text{LN}},\widehat{\pi}_{\text{SNLN}}$ learned by Algorithm 2 and SNLN respectively, under $\delta=0.1$, each with 20 trials.

### v2

@@ 4 Policy Learning under Concept Drift
$\mathbb{E}\big{[}\ell(X,Y(\pi(X));\theta)\,|\,X=x\big{]}$
$=$ $\sum^{M}_{a=1}\mathbbm{1}\{\pi(X)=a\}\cdot\mathbb{E}[\ell(x,Y(a);\theta)\,|\,X=x].$ (8)
@@
$-\mathbb{E}\bigg{[}\ell(X,Y(\pi(X));\bm{\theta}^{*}_{\pi(X)}(X))\bigg{]}.$
@@
Constructing $\widehat{\bm{\theta}}_{\pi}(x)$ with $\widehat{\bm{\theta}}_{\pi(x)}(x)$ substantially reduces the computational complexity of the policy learning task. It is virtually infeasible to estimate $\widehat{\bm{\theta}}_{\pi}(x)$ for each $\pi$ in a policy class $\Pi$ with infinite number of policies. Alternatively, solving for $\widehat{\bm{\theta}}_{\pi(x)}(x)$ transforms the infeasible task of computing a class of infinite nuisance parameters $\{\bm{\theta}_{\pi}(x):\pi\in\Pi\}$ to the feasible task of computing a finite one $\{\bm{\theta}_{a}(x):a\in[M]\}$. It remains an interesting future direction to extend this trick to continuous action spaces.
@@ 4.1 The Learning Algorithm
The policy learning algorithm consists of two main steps: (1) solving for $\bm{\theta}^{*}_{a}$ for each $a\in[M]$ and constructing the policy value estimator $\widehat{\mathcal{V}}_{\delta}(\pi)$; (2) learning the optimal policy $\pi^{*}_{\delta}$ by minimizing $\widehat{\mathcal{V}}_{\delta}(\pi)$.
@@
$G_{a}(x,y):=\ell(x,y;\bm{\theta}^{*}_{a}(x)),~{}\widehat{G}^{(k)}_{a}(x,y):=\ell(x,y;\widehat{\bm{\theta}}^{(k)}_{a}(x)),$
$\bar{g}^{(k)}_{a}(x):=\mathbb{E}\big{[}\widehat{G}^{(k)}_{a}(X,Y(a))\,|\,X=x\big{]}.$
@@
$\widehat{\pi}_{\text{LN}}=\underset{\pi\in\Pi}{\arg\!\max}~{}\widehat{\mathcal{V}}^{\textnormal{LN}}_{\delta}(\pi):=-\frac{1}{K}\sum^{K}_{k=1}\widehat{\mathcal{V}}_{\delta}^{\text{LN},(k)}(\pi);$ (9)
$\widehat{\mathcal{V}}_{\delta}^{\text{LN},(k)}(\pi)=\frac{1}{|\mathcal{D}^{(k)}|}\sum_{i\in\mathcal{D}^{(k)}}\frac{\mathbbm{1}\{A_{i}=\pi(X_{i})\}}{\widehat{\pi}_{0}^{(k)}(A_{i}\,|\,X_{i})}$
$\cdot\big{(}\widehat{G}^{(k)}_{\pi(X_{i})}(X_{i},Y_{i})-\widehat{g}^{(k)}_{\pi(X_{i})}(X_{i})\big{)}+\widehat{g}^{(k)}_{\pi(X_{i})}(X_{i}).$
The above optimization problem can be solved efficiently by first-order optimization methods or policy tree search as in Zhou et al. (2023); we shall elaborate on the implementation in Section 5. The complete policy learning procedure is summarized in Algorithm 2.
@@
  Return: $\widehat{\pi}_{\text{LN}}$ that maximizes $\widehat{\mathcal{V}}_{\delta}^{\text{LN}}(\pi)$ as in Equation (9).
@@ 4.2 Regret Upper Bound
Definition 4.2 (Hamming entropy integral).
@@
The proof of Theorem 4.3 is deferred to Appendix D. At a high level, we decompose the regret and upper bound it with the supremum of the estimation error of policy values:
$\mathcal{R}_{\delta}(\widehat{\pi}_{\text{LN}})=$ $\mathcal{V}_{\delta}(\pi^{\ast})-\mathcal{V}_{\delta}(\widehat{\pi}_{\text{LN}})\leq\mathcal{V}_{\delta}(\pi^{*})-\widehat{\mathcal{V}}_{\delta}^{\text{LN}}(\pi^{*})$
$+\widehat{\mathcal{V}}_{\delta}^{\text{LN}}(\pi^{*})-\widehat{\mathcal{V}}_{\delta}^{\text{LN}}(\widehat{\pi}_{\text{LN}})+\widehat{\mathcal{V}}_{\delta}^{\text{LN}}(\widehat{\pi}_{\text{LN}})-\mathcal{V}_{\delta}(\widehat{\pi}_{\text{LN}})$
$\leq$ $2\sup_{\pi\in\Pi}|\widehat{\mathcal{V}}^{\textnormal{LN}}_{\delta}(\pi)-\mathcal{V}_{\delta}(\pi)|,$
where the last step uses the choice of $\widehat{\pi}_{\textnormal{LN}}$. We bound the above quantity by establishing uniform convergence results for the policy value estimators. Through a careful chaining argument, we show that the dependence of $\mathcal{R}(\widehat{\pi}_{\text{LN}})$ on $n$ is of the order $O(n^{-\frac{1}{2}})$, which is sharper than the $O(n^{-\frac{1}{2}}\log n)$ dependence for that of Mu et al. (2022) by a logarithmic factor. We also note that both regrets are asymptotic in $n$ and hold for sufficiently large $n$.
@@ 4.3 Regret Lower Bound
In this section, we complement the regret upper bound in Theorem 4.3 with a minimax lower bound that characterizes the fundamental difficulty of policy learning under concept drift. Our lower bound is stated in terms of the Natarajan dimension (Natarajan, 1989), defined as follows.
@@
Given an $M$-action policy class $\Pi$, we say a set of $m$ points $\{x_{1},\ldots,x_{m}\}$ is shattered by $\Pi$ if there exist two functions $f_{-1},f_{1}:\{x_{1},\ldots,x_{m}\}\mapsto[M]$ such that $f_{-1}(x_{j})\neq f_{1}(x_{j})$ for all $j\in[m]$ and for any $\sigma\in\{\pm 1\}^{m}$, there exists a policy $\pi\in\Pi$ such that for any $j\in[m]$, $\pi(x_{j})=f_{\sigma_{j}}(x_{j})$. The Natarajan dimension Ndim($\Pi$) of $\Pi$ is the size of the largest set shattered by $\Pi$.
@@
As can be seen from the definition, the Natarajan dimension generalizes the Vapnik-Chervonenkis (VC) dimension (Vapnik & Chervonenkis, 2015) to the multi-class classification setting. The Natarajan dimension is also closely related to the Hamming entropy integral $\kappa(\Pi)$ in our upper bound, as $\kappa(\Pi)=O(\sqrt{\log(d)\text{Ndim}(\Pi)}$(Cai et al., 2020).
@@
Let $\mathcal{P}$ denote the set of all distributions of $(X,A,Y(1),\ldots,Y(M))$ that satisfy Assumption 2.1, 2.4, 3.3, and 3.4.555When we say a distribution $P$ satisfies Assumption 3.4, we mean that under $P$ there exist $\widehat{\bm{\theta}}$, $\widehat{\pi}_{0}$, and $\widehat{g}$ that satisfy the convergence rates in Assumption 3.4. Suppose that $\delta\leq 0.2$, $n\geq\textnormal{Ndim}(\Pi)^{2}$, and $\textnormal{Ndim}(\Pi)\geq 4/(9\varepsilon)$. For any policy leaning algorithm that outputs $\widehat{\pi}$ as a function of $\{(X_{i},A_{i},Y_{i})\}_{i=1}^{n}$, there is $\sup_{P\in\mathcal{P}}\mathbb{E}_{P^{n}}[\mathcal{R}(\widehat{\pi})]\geq\frac{1}{120}\sqrt{\frac{\textnormal{Ndim}(\Pi)}{n\varepsilon}}.$
The proof of Theorem 4.6 is provided in Appendix D.4. Theorem 4.6 implies that for any learning algorithm, there exists a problem instance such that the regret scales as $\Omega(\sqrt{\text{Ndim}(\Pi)/n})$.
@@
Recalling the relationship between the Natarajan dimension and the Hamming entropy integral in the remark above, we see that our proposed algorithm achieves the minimax rate in the sample size and the policy class complexity.
Metric $\delta$ Policy $n=$7500 $n=$13500 $n=$16500 $n=$19500
$\bar{\mathcal{V}}_{\delta}$ 0.05 $\widehat{\pi}_{\text{LN}}$ 0.2272±$\pm 0.002$ 0.2299±$\pm 0.001$ 0.2303±$\pm 0.001$ 0.2310±$\pm 0.001$
@@
Table 2: Empirical robust policy value $\bar{\mathcal{V}}_{\delta}$ of policies $\widehat{\pi}_{\text{LN}},\widehat{\pi}_{\text{SNLN}}$ on the simulated dataset and the corresponding, over 50 seeds.
Figure 1: Empirical robust policy value $\bar{\mathcal{V}}_{\delta}$ of policies $\widehat{\pi}_{\text{LN}},\widehat{\pi}_{\text{SNLN}}$ on the real-world dataset, over 50 seeds. Shading corresponds to $95\%$ confidence intervals.
Metric Policy $n=$7500 $n=$13500 $n=$16500 $n=$19500
@@
Table 3: Empirical worst case policy reward on the KL-sphere $\tilde{\mathcal{V}}_{\delta}^{\min}$ of policies $\widehat{\pi}_{\text{LN}},\widehat{\pi}_{\text{SNLN}}$, over 20 seeds.
5 Numerical Results
We evaluated our learning algorithm in two settings: a simulated and a real-world dataset, against the benchmark algorithm SNLN in Si et al. (2023, Algorithm 2).
Simulated Dataset. Our data generating process follows that of the linear boundary example in Si et al. (2023). We let the context set $\mathcal{X}$ to be the closed unit ball of $\mathbb{R}^{5}$ and let the action set to be $\{1,2,3\}$; the rewards $Y(a)$’s are mutually independent conditioned on $X$ with $Y(a)\mid X$ being Gaussian, for $a\in[3]$. The training datasets $\mathcal{D}_{\text{train}}$ are generated with an unknown given behavior policy $\pi_{0}$ over 50 random seeds. We similarly generate 100 testing datasets $\mathcal{D}_{\text{test}}$ of size 10,000. The details are given in Appendix C.
Real-world Dataset. We consider the dataset of a large-scale randomized experiment comparing assistance programs offered to French unemployed individuals provided in Behaghel et al. (2014). The decision maker is trying to learn a personalized policy that decides whether to provide: (i) an intensive counseling program run by a public agency ($A=0$); or (ii) a similar program run by private agencies $(A=1)$, to an unemployed individual. The reward $Y$ is binary and indicates reemployment within six months. The processed dataset is provided in Kallus (2023).
Implementation. In our implementation, the number of splits is taken to be $K=3$. We use the Random Forest regressor from the scikit-learn Python library to estimate $\widehat{\pi}_{0}$ and $\widehat{g}$. For estimating $\bm{\theta}^{*}$, we adopt the cubic spline method and employ the Nelder-Mead optimization method in SciPy Python library (Virtanen et al., 2020) to optimize the coefficients in the spline approximation, where the obtained estimator has threshold at 0.001 to guarantee Proposition 2.5. Finally, we optimize and find $\widehat{\pi}_{\textnormal{LN}}$ with policytree (Sverdrup et al., 2020). [^6] A working example on the real-world dataset is given in https://github.com/off-policy-learning/concept-drift-robust-learning.
The benchmark algorithm SNLN is adapted from Si et al. (2023, Algorithm 2) as in Kallus et al. (2022). [^7] The implemented code of SNLN benchmark is provided by Kallus et al. (2022). Since Si et al. (2023, Algorithm 2) is designed for joint distribution shift formulation, we revised the original algorithm to fit our concept drift setting. The well-known KL-divergence chain rule (Cover, 1999) gives
$D_{\text{KL}}(Q_{X,Y}\,\|\,P_{X,Y})$
$=D_{\text{KL}}(Q_{X}\,\|\,P_{X})+D_{\text{KL}}(Q_{Y\,|\,X}\,\|\,P_{Y\,|\,X}).$ (10)
Therefore, given any uncertainty set radius $\delta$ and known covariate shift (in this experiment, we assume no covariate shift), Si et al. (2023, Algorithm 2) can be used to implement policy learning under concept drift. Note that SNLN admits known propensity scores. As we only consider the case where the propensity scores are unknown, we complement Si et al. (2023, Algorithm 2) with estimated propensity scores from Random Forest Regressor in scikit-learn, the same way as in the implementation of Algorithm 2.
Evaluation. For a learned policy $\widehat{\pi}$, we evaluate its performance by the following performance metrics. (i) We use the testing dataset $\mathcal{D}_{\text{test}}$ to estimate the robust policy value $\mathcal{V}^{*}_{\delta}(\widehat{\pi})$ by the empirical robust policy value
$\bar{\mathcal{V}}_{\delta}(\widehat{\pi})=-\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{i\in\mathcal{D}_{\text{test}}}\ell(X_{i},Y_{i}(\pi(X_{1}));\bm{\theta}_{\widehat{\pi}}(X_{i})),$
where the nuisance parameters $\bm{\theta}_{\widehat{\pi}}(X)$ are found via cubic spline method and employ the Nelder-Mead optimization method using the testing dataset according to policy $\widehat{\pi}$: $\mathcal{D}_{\text{test},\widehat{\pi}}=\{(X_{i},Y_{i}(\widehat{\pi}(X_{i})))\}$. (ii) We also perturb the simulated dataset to mimic possible real-world distributional shift. For each testing dataset $j$ containing 10000 data points of the total 100 testing datasets, we generate a new testing dataset, with each reward distribution $(\tilde{Y}_{i}^{(j)}(1),\tilde{Y}_{i}^{(j)}(2),\tilde{Y}_{i}^{(j)}(3))$ randomly sampled on the KL-sphere centered at the reward distribution $(Y_{i}^{(j)}(1),Y_{i}^{(j)}(2),Y_{i}^{(j)}(3))$ of the testing data point with a radius $\delta$. Then we evaluate $\widehat{\pi}$ using
$\tilde{\mathcal{V}}_{\delta}^{\min}(\widehat{\pi}):=\min_{j\in[100]}\bigg{\{}\frac{1}{10000}\sum_{i=1}^{10000}\tilde{Y}_{i}^{(j)}\big{(}\widehat{\pi}(X_{i}^{(j)})\big{)}\bigg{\}}.$
Results. Table 2 and 1 reports the values $\bar{\mathcal{V}}_{\delta}$ of the learned policies $\widehat{\pi}_{\text{LN}}$ and $\widehat{\pi}_{\text{SNLN}}$ on the simulated and the real-wrold dataset, respectively. Table 3 provides the result of $\tilde{\mathcal{V}}_{\delta}^{\min}$. All results are reported with 95% confidence intervals. Table 2 shows that $\widehat{\pi}_{\text{LN}}$ outperforms the benchmark $\widehat{\pi}_{\text{SNLN}}$ consistently, with higher policy values and similar 95% confidence intervals. In Figure 1, $\widehat{\pi}_{\text{LN}}$ continues to show this advantage over $\widehat{\pi}_{\text{SNLN}}$ on the real-world dataset. With a higher $\delta$, the policy values of $\widehat{\pi}_{\text{LN}},\widehat{\pi}_{\text{SNLN}}$ are smaller, due to a bigger uncertainty set. Table 3 shows that $\widehat{\pi}_{\text{LN}}$ achieves higher worst-case rewards than $\widehat{\pi}_{\text{SNLN}}$ does, in a more realistic setting with concept drift testing datasets. Together, we see that $\widehat{\pi}_{\text{LN}}$ succeeds in finding a better policy under concept drift; while the performance of $\widehat{\pi}_{\text{SNLN}}$ is comprised by its conservative policy learning process, in which it considers joint distributional shifts even though it is given the information that no covariate shifts took place.
The results align with the intuition that Algorithm 2 admits a subset of the uncertainty set that SNLN considers, as explained in Equation (5). Consequently, $\mathcal{V}_{\delta}(\widehat{\pi}_{\text{SNLN}})$ is a lower bound of $\mathcal{V}_{\delta}(\widehat{\pi}_{\text{LN}})$ in theory, and by the results in Table 2, in practice. In real-world applications, knowing the source of the distribution shift effectively shrinks the uncertainty set, thereby yielding less conservative results. Since it is fairly easy to identify covariate shifts (comparing to detecting concept drift), when the decision maker observes none or little covariate shifts and would like to hedge against the risk of concept drift, it is suitable to apply our method which outperforms existing method designed for learning under joint distributional shifts.
One limitation of our methodology (as well as in other DRO works) is the choice of $\delta$. The parameter $\delta$ controls the size of the uncertainty set considered and thus controls the degree of robustness in our model — the larger $\delta$, the more robust the output. The empirical performance of the algorithm substantially depends on the selection of $\delta$. A small $\delta$ leads to negligible robustification effect and the algorithm would learn an over-aggressive policy; a large $\delta$ tends to yield more conservative results. A more detailed discussion can be found in Si et al. (2023).
In Appendix C, we also provide simulation results of Algorithm 1 for a fixed target policy, which show that Algorithm 1 can estimate the distributionally robust policy value under concept drift efficiently.